Solve (type an integer or a simplified fraction. Use a common to separate answers as needed. Type each solution only once)
w^2+w-20=0

Answer by mananth(16946)

w^2+w-20=0
...
w^2+5w-4w-20
w(w+5)-4(w+5)=0
(w+5)(w-4)=0
...
w= -5, 4